OpenStudy (anonymous):

Help with Algebra 1 Rational Expressions?!

OpenStudy (anonymous):

What is the simplified form of x^4-81 over x+3 A) x^3-3x^2+9x-27 B) (x^2+9)(x^2-9) C) x^3+3x^2+9x+27 D) x^3+3x^2-9x-27 2)What is the simplified form of (x^2yz)^2(xy^2z^2) over (xyz)^2? A) (xyz)^2(xyz) over x B) 1 over (xyz)^2 C) x^3y^2z^3 D) x^4y^4z 3) What is the excluded value of the rational expression 2x+6 over 4x-8? A) -3 B) -2 C) 0 D) 2 4) What is the simplified form of x^2-25 over x^2-3x-10 A) x-3 over x-5 B) (x-5)(x+5) over (x+3)(x-3) C) x+5 over x+2 D) x-5 over x+2

OpenStudy (unklerhaukus):

\[x^4-81 \over x+3 \\ =\frac{(x^2+9)(x^2-9)}{x+3}\\ =\frac{(x^2+9)(x+3)(x-3)}{x+3}\\ =\]